William is 2 times as old as Stephanie. 28 years ago, William was 9 times as old as Stephanie. how old is Stephanie now?
Answer by stanbon(75887)

William is 2 times as old as Stephanie. 28 years ago, William was 9 times as old as Stephanie. how old is Stephanie now?
----
Equations:
W = 2*S
-----------
W-28 = 9(S-28)
----
Substitute for "W" and solve for "S":
2S-28 = 9S-9*28
-----
7S = 8*28
S = 8*4
Ans: S = 32 yrs (Stephanie's age now)
-----------
Cheers,
Stan H.
-----------